Suppose the demand equation in the market for packages herbal tea is D=140-4P and the supply equation in the market for packages of herbal tea is S=60, calculate and graph the equation.

I got MCP = $20, and MCQ = 60 packages. However, I don't remember on which axis to graph the supply equation.
Answer #1

Demand is Q = 140 - 4P and supply is Q = 60. Quantity demanded and supplied (Q) is drawn on horizontal axis and price (P) is drawn on vertical axis. Here at the equilibrium we have 140 - 4P = 60 or P = (140 - 60)/4 = 20 and quantity = 60 units (fixed). Graph is shown below

Demand and supply 50 E 20 10 0 10 20 40 50 60 70 80 90 100 110 120 130 140 Quantity - Supply Demand 30 40 30 Price